OpenStudy (anonymous):

integral of (1/((2x-1)^1/3)) dx

OpenStudy (anonymous):

\(u=2x-1, du =2dx,\frac{1}{2}du = dx\) then integrate \[\frac{1}{2}\int u^{-\frac{1}{3}}du\]

OpenStudy (anonymous):

so would it be 1/2 ln| (2x-1)^1/3| + c

OpenStudy (anonymous):

?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

there is no log here, use the power rule backwards

OpenStudy (anonymous):

\[\int u^{-\frac{1}{3}}du=\frac{3}{2}u^{\frac{2}{3}}\]

OpenStudy (anonymous):

So then its 3/4 (2x-1)^2/3

OpenStudy (anonymous):

?